Please help me solve this equation: 3(2x-1)-11= -2 (1-x)
I thought I had done it right a number of times, but when I plugged in the number I said was equal to x, it was completely wrong! All I know for sure is the Distributive Property has to be done first.

Answer by rfer(16322)

3(2x-1)-11=-2(1-x)
6x-3-11=-2+2x
4x=12
x=3